# NMCSE: Noise-Robust Multi-Modal Coupling Signal Estimation Method via Optimal Transport for Cardiovascular Disease Detection

### v1

@@ 5.2.1 Experimental Setup
    Used clean ECG and PCG recordings from PhysioNet 2016 dataset as baseline
@@ 5.2.2 Results and Analysis
5.2.3 Parameter Sensitivity Analysis
@@
Table 3 shows that optimal performance is achieved at $\alpha=1.0$ and $\beta=0.1$, which is consistent with our theoretical analysis. Insufficient amplitude weighting ($\alpha<0.5$) leads to degraded performance with MSE increasing by up to 63% at $\alpha=0.01$. Similarly, performance degrades when temporal weighting ($\beta$) exceeds 0.1, with a 9% increase in MSE at $\beta=1.0$. Within the tested parameter range, NMCSE shows moderate sensitivity to parameter selection, with performance variations of less than 10% when parameters remain within 50% of their optimal values. The optimal $\alpha/\beta$ ratio remained consistent across different noise levels (5-30 dB), suggesting that once calibrated, the same parameter configuration can be used across varying clinical conditions.
@@ 5.3.1 Ablation Experiment
As shown in Table 4, among single-modal inputs, the NMCSE-estimated coupling signal achieved the highest performance (90.25% accuracy, 0.93 AUC), outperforming both ECG (81.23%) and PCG (84.71
@@
5.3.2 Comparison with Existing Multi-Modal Methods

### v2

@@ 5.2.1 Experimental Setup
    Used clean ECG and PCG recordings from PhysioNet dataset as baseline
@@ 5.2.2 Results and Analysis
5.2.3 Spectral Coherence Analysis
To understand the frequency-domain advantages of NMCSE, we performed spectral coherence analysis comparing reference coupling signals with those estimated by different methods. Fig. 4A shows that NMCSE maintains significantly higher coherence with the reference signal in the mid-frequency range (10-100 Hz), with approximately 31% higher coherence compared to deconvolution. This frequency band contains critical mechanical cardiac events, including the first (S1) and second (S2) heart sounds.
The coherence difference map (Fig. 4B) demonstrates how NMCSE’s advantage persists across varying noise levels, with the most pronounced benefits in the mid-frequency band. As SNR decreases, NMCSE’s robustness becomes increasingly valuable, particularly at lower SNR levels (5-10 dB) typical in clinical environments.
Figure 4: Spectral coherence analysis of coupling signal estimation methods. (A) Magnitude squared coherence between reference and estimated coupling signals (10dB SNR), highlighting NMCSE’s superior preservation of mid-frequency components. (B) Coherence difference map across frequencies and noise levels. (C) Average coherence preservation by frequency band (10dB SNR), quantifying NMCSE’s 28% improvement in mid-frequency preservation.
Quantitative analysis (Fig. 4C) confirms that NMCSE preserves 28% more coherence in the mid-frequency band compared to deconvolution at 10dB SNR. This explains why NMCSE-estimated coupling signals retain more diagnostically relevant features. At 10dB SNR, NMCSE preserves 87% of clinically relevant features compared to 73%, 68%, and 64% for sparsity-based, Wiener, and Tikhonov methods, respectively.
The frequency-specific performance improvement demonstrated by NMCSE aligns with our theoretical analysis in Section 3.7, confirming that our distribution-matching approach effectively mitigates noise amplification across the frequency spectrum, with particularly strong benefits in physiologically significant frequency bands.
@@ 5.2.3 Spectral Coherence Analysis
5.2.4 Parameter Sensitivity Analysis
Table 3 shows that optimal performance is achieved at $\alpha=1.0$ and $\beta=0.1$, which is consistent with our theoretical analysis. Insufficient amplitude weighting ($\alpha<0.5$) leads to degraded performance with MSE increasing by up to 63% at $\alpha=0.01$. Similarly, performance degrades when temporal weighting ($\beta$) exceeds 0.1, with a 9% increase in MSE at $\beta=1.0$. Within the tested parameter range, NMCSE shows moderate sensitivity to parameter selection, with performance variations of less than 10% when parameters remain within 50% of their optimal values. The optimal $\alpha/\beta$ ratio remained consistent across different noise levels, suggesting that once calibrated, the same parameter configuration can be used across varying clinical conditions.
@@ 5.3.1 Ablation Experiment
As shown in Table 4, among single-modal inputs, the NMCSE-estimated coupling signal achieved the highest performance (90.25% accuracy, 0.93 AUC), outperforming both ECG (81.23%) and PCG (84.71%) alone. This demonstrates that the coupling signal effectively captures the cross-modal relationship between electrical and mechanical cardiac activities.
@@
5.3.2 Comparison with Existing Multi-Modal Methods